Home‎ > ‎Reports‎ > ‎Crew104‎ > ‎

day08


Crew 104 Log Book for April 30, 2011

Photos of the day





Commander's Check-In Report

Jon Clarke Reporting

Crew Physical Status:  Good but tired

Time Departed/Returned from EVA:  Departed at 9:00 for the drill site, returned at 12:00 for lunch.  Back on site by 2:00, and back indoors by 5:00

Brief Narrative of Field Mission Results:  The MARTE drill is non-operational because of deformation of parts and cannot be repaired on site.  So a Miluwalkee core drill was used instead as a backup.  In three hours it drilled as much core as the MARTE drill did in two days.  Disadvantages of the method include contamination by water of which it uses a lot to cool the bit and remove the cuttings, and the amount of power required.  But it is definitely the type of drill a human crew would require on the Moon or Mars.  A challenge for the engineers to come up with waterless drilling.

EVA Data/Interpretations:  Analysis of drill performance, XRD mineralogy, and sample DNA and organics are pending.

Engineering/Hab Maintenance: No outstanding issues

Report Transmission Schedule:  Commander's and engineering reports pending.

Plans for Tomorrow: Probably further analysis of samples.  Being Sunday and the rotation midpoint, we will probably have a fairly quiet day.

Inventory: no items outstanding

Support Requested: None

Miscellaneous:  Not being in sim we have two most welcome visitors, Shannon and Daniel.  They will camp for the weekend.

Commander's Report

Jon Clarke Reporting

Quite a day today. The weather was cold with flakes of snow. But there was no wind, which was an improvement.

The MARTE drill was declared off-line for the duration of the rotation, but despite this we achieved our primary mission goals with this system. To continue researching drilling, we substituted an alternative drill, a Milwaukee drill that is designed to collect short cores. This was highly successful, fast and efficient. It will be suitable for geological and organic geochemical drilling, but probably not for detection of microorganisms because of contamination from the drilling water. Such a drill would be ideal for use on the Moon and Mars, provided an alternative drilling fluid could be used.

Those not drilling spent a lot of time running organic and DNA and XRD analyses. ATV training was also undertaken for those who still had not had it.

Lastly we have visitors, Shannon Rupert and Daniel. Over dinner Shannon entertained us with stories of MDRS and FMARS missions of yore and there was much discussion about the present and future of the facility.

We are all tired tonight and are looking forward to a slightly quieter day tomorrow. Our plans are weather dependent, they include looking for more different lithology to drill, especially limestones, further lab analyses, and some infill sampling.


Engineering Report

David Willson Reporting

Generator/Electricity:

Kitty (Used? Yes/No): Used Continuously and not stopped

Honey (Diesel Generator) (Used? Yes/No): Not Used

Wendy (Gas Generator) (Used? Yes/No): Not Used

Battery Charge Level: Not Checked

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Generator operating very well continuously and not stopped

Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 24% in Tank

Propane Tank (Full/0.75/0.5/0.25/Empty): 30%

Gasoline Tank (Full/0.75/0.5/0.25/Empty): N/A

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): Full

Honey Oil Quantity (# of Quarts in storage): Still Very Low, Will be toped up tomorrow

ATV Oil Quantity (# of Quarts in storage): Not Checked

Car Oil Quantity (# of Quarts in storage): Not Checked


Water Status:

Outside Potable Water Tank Level (inches from bottom): 75% Full

Trailer Potable Water Tank Level (inches from bottom): 15 cm

Hab Potable Water Tank Level (inches from bottom): 44%

Potable Water Meter Reading: 458769

In to GreenHab Meter Reading: N/A

Notes/Comments: Green Hab not operational but water is being recycled


GreenHab:

Condition of plants in Tank 1: 5%

Condition of plants in Tank 2: Full

Condition of plants in Tank 3: 15% Full

Greenhab Notes/Comments: Green house not used


Observatory:

Telescope Used? (Yes/No): NO

Observatory Notes/Comments:


Transportation:

Hab Car (NOT RENTAL): Operating

Used (yes/no): NO

Oil Checked (yes/no):

Tire Status:

Odometer at end of day:

Notes/Comments on Hab Car:


ATV 1 (Opportunity): Operating

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


ATV 2 (Spirit):

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


ATV 3 (Viking I):

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


ATV 4 (Viking II):

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): A/C not operating 70 degrees F

Actual temperature upstairs (Farenheit): 68 degrees F

Thermostat setting downstairs (Farenheit): 63 degrees F

Actual temperature downstairs (Farenheit):


Computers/Networking Infrastructure: Operating

Notes/comments on internet/computers?: All OK


General Engineering Notes/Comments:

Photo of missing connection on ATV Viking 1.

New connection has been purchased but not yet installed.